OpenStudy (anonymous):

The function f(x) = 245(4)x represents the growth of a fruit fly population every year in the orange grove. Stacey wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Stacey's purposes? @ranga

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Sorry, I'm drawing a blank at the moment.

OpenStudy (kc_kennylau):

you mean \(\large254\times4^x\)?

OpenStudy (anonymous):

yes @kc_kennylau

OpenStudy (kc_kennylau):

This is the number of flies itself, or the change in the number of flies?

OpenStudy (anonymous):

the growth

OpenStudy (kc_kennylau):

so just 12-root it lol?

OpenStudy (anonymous):

@kc_kennylau how do i do that

OpenStudy (kc_kennylau):

\[\sqrt[12]{254\times4^x}\]

OpenStudy (anonymous):

uhh how do you solve that @kc_kennylau

OpenStudy (kc_kennylau):

lolz sry idk :/

OpenStudy (anonymous):

@kc_kennylau huh you dont know how to solve it?

OpenStudy (kc_kennylau):

exactly xD

OpenStudy (anonymous):

@ganeshie8 pls help

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@rm152

OpenStudy (ranga):

In the formula x is in years. If x should be in months, then replace x in the formula with x/12 and now x will be in months.

OpenStudy (anonymous):

so how would it look

OpenStudy (ranga):

\[f(x) = 245*4^{x/12}\]

OpenStudy (anonymous):

\[f(x)=254*4^(x/12)\]

OpenStudy (anonymous):

so what do i plug in?

OpenStudy (ranga):

Are there answer choices?

OpenStudy (anonymous):

You plug in the number of months that have elapsed.

OpenStudy (anonymous):

@ranga thanks ranga and rm152

OpenStudy (anonymous):

ganga can you give a medal to rm152

OpenStudy (ranga):

sure.

OpenStudy (anonymous):

thanks:)

OpenStudy (ranga):

You are welcome.